OpenStudy (anonymous):

Leif started making contributions to a Traditional IRA when he got his first job, at age 32. His contributions averaged $3,200 annually. Leif was in the 20% tax bracket during his working years, when he continued to make these annual contributions. The average annual rate of return on the account was 6.8%. Upon retirement at age 69, Leif stopped working and making IRA contributions. Instead, he started living on withdrawals from the retirement account. At this point, Leif dropped into the 10% tax bracket. Factoring in taxes, what is the effective value of Leif’s Traditional IRA at retirement?

OpenStudy (anonymous):

Assume annual compounding Answers: $440,728.02 $305,314.21 $483.080.97 $391,758.24

OpenStudy (tkhunny):

69-32 = 37 ==> 3200 * 1.068^37 69-33 = 36 ==> 3200 * 1.068^36 69-34 = 35 ==> 3200 * 1.068^35 69-35 = 34 ==> 3200 * 1.068^34 etc. 69-67 = 2 ==> 3200 * 1.068^2 69-68 = 1 ==> 3200 * 1.068^1 Where is this leading?

OpenStudy (anonymous):

69-69=0 ==> 3200 * 1.068^0

OpenStudy (tkhunny):

Okay, now add them all up? 3200 * 1.068^37 + 3200 * 1.068^36 + ... + 3200 * 1.068^0 3200 * (1.068^37 + 1.068^36 + ... + 1.068^0) \(3200\cdot \dfrac{1.068^{38} - 1.068^{1}}{1.068 - 1}\) Now what>

OpenStudy (anonymous):

Thank you!